Using a quadratic equation (if possible) please solve the following:
The sum of 2 numbers is 48 and their product is 432. Find the numbers.

Answer by richwmiller(17219)

x+y=48
x*y=432
y=48-x
x*(48-x)=432
48x-x^2-432=0
-(x-36) (x-12) = 0
x=12
y=36